# GRAPH - An readout ASIC for large MCP based detectors

### v1

@@ 1 Introduction
In view toward space grade environmental specifications, aside the radiation hardness which are mission driven, and not specifically addressed here. The most concerning is the power consumption, as heat being very difficult to dissipate, and the system’s signal to noise ratio capable to bring the desired pixel resolution smaller than 20 $u$m FWHM. at low MCP bias voltages. Our previous work brought us to the conclusion, that the following parameters should be met in view of designing a new ASIC, and is summarized in table 1.
@@
Event readout rate 1 $M$event/s/ASIC (1 event = 48 samples)
@@
Space missions differ and in the technology selection there is no generic solution, but the choice falls on matching trade offs on many levels, to fit mission’s purpose. In this respect, we note that there are at least two groups working on a similar endeavor. An implementation using a pixelated pattern, opposed to PXS, was under study using the Timepix [9], with an updated version with the Timepix4 ASIC [10], [11]. The other group uses the Beetle ASIC [12] to accomplish a similar endeavor using the PXS approach.
@@ 2.1 Architecture overview
Redesigning into an integrated system having amplifiers and ADCs required moving to a smaller node (250nm to 130nm), which has some technical constrains. A waveform sampling array is made of a large number of switches and capacitors, which begin to channel leak charge from the storage capacitors back to the input, hence loosing the correct sampled value. This effect was found in simulation to be in the range of a few mV/us. Despite short access and conversion times, opting for a random digitization moment of the event in the sampling array would have non-deterministic effects, hence produce large conversion errors. In order to be able to retain a high S/N ratio of the system, the time between sampling and conversion should remain constant for all chips running synchronously. We explore, to our knowledge, a somewhat novel ADC architectural concept using a mixed signal memory. We named it the Hybrid Universal sampLing Architecture (HULA).
@@
The front end is composed of programmable CSAv3 [4] amplifiers with an additional circuit that allows bypassing the CSA, if the ASIC is used in another applications, and the amplified signal is then fed into the HULA ADC. In order to control the sampling and conversion, the Timebase circuit generates the sampling vector address from a single input clock pin. It also generates internal synchronization signals that govern the conversion. In addition, an independent read system mechanism is available, which minimizes the readout dead time for this particular case of use. The slow control system is composed of a 4 wire SPI like receiver, accepring 1.2V LVCMOS signals. Its command decoder accepts frames being 64 bit long, containing the targeting internal address space and values to control circuits for various settings and internal biasing voltages (DACs). Figure  2(a) depicts the internal structure of the ASIC and, Fig 2b the pictogram of the chip.
With the GRAPH ASIC redesign we address also some desired improvements on the system level, and these are:
@@ 2.2 Sampling windows and memory array
A single sample of the HULA architecture is built using a mixed signal memory, employing a small 30 fC capacitor for analog sampling, a comparator, and a 12 bit digital memory SRAM alike, shown in figure  3(a). Once the sample acquires its analog value, a conversion using the Wilkinson ADC approach will take place at some deterministic point in time. While in conversion, the 12 SRAM bits are connected to the bus and follow the counter value. As soon as the ramp value is above the signal in the capacitor, a D flip flop is set, which disconnects the 12 bit memory from the bus, retaining the last counter value. A separate bus is used to independently read the memory array.
A single sample is stacked into a so called vertically oriented window containing 64 samples, and there are two banks having 16 windows for each channel. Therefore the memory length of 2048 samples is available for every channel. The memory banks are stacked horizontally, while channels are stacked vertically 3b.
@@ 2.3 Timebase generation circuit
The Timebase generation circuit, figure 4(a) is composed of a self looped serial shift register being 64 bits in length, inside which a sample actuation signal of programmable width is let to circulate in a loop. When the external nCLR signal is released, the shift of this signal occurs driving the sampling action through the samples in the window at an externally provided clock. At each turnaround, a 5 bit counter increments by one, and its value represents the window which is addressed for sampling, figure  4b. The signal from the shift register is directly fanned out to all windows arrays, while the decoded value in the 5 bit counter selects the window undergoing sampling  3(a). In addition, the Timebase extrapolates a few more signals necessary for internal synchronization. An AnB signal is sent off the the chip, when the transition from one bank to another happens and internal signals are set to control the conversion controller.
Figure 4: (a)Functional block diagram of the Timebase circuit (b) Sampling signal sequence diagram
@@ 2.4 Conversion controller
The main function of the conversion controller is to execute the conversion at a deterministic moment. It is composed of ramp generator, which relies on a programmable current source driven into a capacitor to create a voltage ramp  5(a). A gray coded counter value is derived from a synchronous counter that revolves at the externally provided frequency clock. The 12 bit Gray code and analog ramp signal are further buffered for fanout over the 16 channels of the entire bank  5b. Hence the ASIC has two such controllers, one for each bank.
@@ 2.5 HULA operation
HULA operates on a continuously revolving pattern. Once the sampling starts, the analog signal is sampled in the first Bank. As soon as the first Bank A has completed sampling, the AnB signal changes, signaling the transition to the external world, and automatically starts to sample bank B and puts Bank A into conversion. When the Bank B has completed sampling, Bank A contains digitized data from the previous revolution which are available for the duration of new sampling period of Bank A, figure  6 .
Considering there are 1024 samples in each Bank, selecting a sampling clock of 125 MHz will require 8,192 $\mu$s to complete the sampling of one Bank. To reach 12 bit resolution in this time, the external clock shall run at least at 2 $n$s period, hence at 500 $M$Hz, and the ramp should be adjusted to cover the operating headroom. This allows to set an arbitrary sampling frequency, if a lower sampling rate is adequate. It also enables reading a much larger portion of the memory in the given conversion time slot, covering more events.
@@ 2.6 Data readout
The mechanism for memory data readout is completely independent from sampling and conversion, and allows some flexibility in accessing the desired address. It is possible to read consecutive samples from the chosen address, while loading a new one, reducing dead time. This is achieved having a receiver composed of 3 serial inputs (channel,window, sample), a corresponding read clock and load confirmation. When an address is shifted in the system and loaded, the value is an offset added to a counter, which is driven by the read clock. At the rising edge of a clock cycle a new vector is decoded pointing to a single cell. This gets attached to the read bus and streamed to a 12 bit parallel output of the chip, which an FPGA latches at the opposite clock edge, figure  7a. Introducing a value of zero to all three registers, and continuing clocking, one would read sample 0 from channel 0, till the last sample of the channel 0. By continuing asserting the read clock, sample 0 of channel 1 would follow and so on. In our application, we load the address of interest, and make the next 6 clock cycles to load new address, while collecting requested data from the previous call. Such access is envisioned for pulse shape extraction. The maximum transmission frequency tested so far was 60 MHz, which would drain the entire memory in loop mode in about 0.5 $m$s, collecting some 50kB of data.
@@ 2.7 Trigger section
Unlike operating in loop mode, in continuous mode triggers are key for the selection of events to pull from GRAPH memory. Each channel has a trigger circuit composed of a programmable threshold, a logical selection for firing on rising or falling edge of the input signal, and a self reset logical signal (the actual trigger output), which pulse width can be set between a ns to few hundred ns by an internal DAC, figure 8(a). Adjacent triggers are OR-ed and fanned out toward the FPGA via an LVDS output 8b. To a degree it allows to discriminate the pulse hit, by setting a common threshold, while setting the width differently. This configuration notably simplifies the required firmware. Ideally, the FPGA is only required to have an 11 bit counter (2048 sample space) being released by the nCLR signal as for all chips, and increments synchronously with the sampling clock. At trigger arrival the value is latched, and the readout processing sequence can estimate the events of interest to extrapolate the charge cloud signal form the MCP/anode. An additional possibility to enhance the functionality is by adding a TDC on the muxed triggers, and locate with better timming the absolute time of the photon arrival. By testing the system a sub $n$s time of arrival precision should be possible. An additional option, the software trigger, is also available, and can be initiated via the slow control serial register. Both trigger options would stop the ASIC in loop mode operation, as well as, both can be masked to be overruled.
@@ 3 Results
For testing the ASICs, a PCB board that can be attached to a detector was produced. This solution has currently 4 ASICs wire bonded directly onto the PCB along with a minimal set of low drop regulators and a subsets of dual inline connectors to connect to an external FPGA board. For testing, the sampling clock was set to 125MHz, and the Wilkinson conversion clock was set to 500 MHz. Initially, the CSA amplifier was skipped, and only the performance of the HULA core was examined by carefully tuning the conversion slope to match the pedestal offset on both Banks. Because the comparators in each memory cell cannot operate near the voltage supply rails, using a DC swap the usable ADC headroom was extracted. Although each ASIC is slightly different, we found all of them operate safely within a 800 mV headroom covered by some 3000 ADC counts. The INL has found to be of few percent, figures 9a,9b.
By applying a DC voltage of about 600 mV to the input, the recorded pedestal spread was found to be considerable, with a sigma of 142 counts, figure 10a This somewhat expected result comes from a combination of factors. These are the systematic production mismatch of the capacitors used in the sampling cell, the intrinsic firing point error of the comparator, and the location of cell itself inside the large memory array, as the ramp propagates through the fanout. A solid calibration method is to measure each cell about a 100 times and average its value. As a new measurement is taken, the averaged offset value is subtracted from the measurement. In doing so, the pedestal value is removed and the intrinsic offset of the cell is corrected reducing the spread to a sigma of 23 counts, figure 10b. The ADC has an SNR of 20log*(3000/23) of around 42dB, with the sensitivity of about 6 mV. Although the system is set to use 12 bits of data the ENOB figure is more around 7.
@@
The system was connected to the detector, and the CSAv3 amplifiers were turned on having the calibration procedure being repeated. The amplifier noise generated by the 5 $p$F input capacitance has worsen the result to 38 ADC counts, figure 11a. Some of that was found that could be reduced by off line post processing the data using a combination of a median and FIR filter to some degree of success, figure 11b. None the less, the obtained raw results are very close to the specifications and expected performance.
@@ 4 Summary and future work
We developed the GRAPH ASIC, which enables the construction of a readout system intended to couple to a large area MCP single photon imaging detector, designated for future space applications. The test board was constructed and the ASIC performance evaluated. Furthermore initial test confirm the correct functionality as well as correct operation. In addition tests recording detector signals were carried out. Considering the low power consumption of 50mW per channel, an extrapolated value for a 8x8 $c$m detector requiring 128 channels per axis *(16 ASICs) would yield a power consumption of 13 Watts FPGA excluded. In the future we look to test a new kind of firmware, that would only record events, by asserting a few clock cycles only when a signal is detected via triggers, and force conversion, if an arbitrary overdue time has passed to avoid losing the signal due to leakage. In this way, we aim to optimize the usage of the memory while preserving the sampling clock. Ultimately, an improbev version of HULA with a much better SNR could operate as an event sampler, digitize and data hold system.

### v2

@@ 1 Introduction
In view toward space grade environmental specifications, aside from radiation hardness which is mission-driven, and not specifically addressed here, the most concerning is the power consumption. Heat is very difficult to dissipate, and the system’s signal-to-noise ratio and sensitivity must be able to bring the desired pixel resolution toward 20 $u$m FWHM at low MCP bias voltages. Our previous work brought us to the conclusion, that the following parameters should be met in view of designing a new ASIC, and is summarized in table 1.
@@
Event readout rate 1 M event/s/ASIC (1 event = 6 samples*8 channels)
@@
Space missions differ and the technology selection doesn’t offer a generic solution, trade-offs are made to fit the mission’s purpose. In this respect, we note that there are at least two groups working on MCP based high resolution single photon imaging detectors. An implementation using a pixelated pattern, opposed to PXS, was under study using the Timepix [9], with an updated version with the Timepix4 ASIC [10], [11]. Another group uses the Beetle ASIC [12] to accomplish a similar endeavor using the PXS approach.
@@ 2.1 Architecture overview
Redesigning into an integrated system with amplifiers and ADCs required moving to a smaller node (250 $n$m to 130 $n$m), which comes with some technical constrains. A waveform sampling array is made of a large number of switches and capacitors. These begin to channel leak charge from the storage capacitors back to the input, hence loosing the correct sampled value. This effect was found in simulation to be in the range of a few mV/us. Despite relying on short access and conversion times, opting for a random digitization moment of the event in the sampling array would have non-deterministic effects, hence produce large conversion errors. In order to be able to retain a high S/N ratio of the system, the time between sampling and conversion should remain constant for all chips running synchronously. We explore, to our knowledge, a somewhat novel ADC architectural concept using a mixed signal memory. We named it the Hybrid Universal sampLing Architecture (HULA).
@@
The front end is composed of programmable CSAv3 [4] amplifiers with an additional circuit that allows bypassing the CSA, if the ASIC is used in applications, where the amplified signal is fed into the HULA ADC. In order to control the sampling and conversion, the Timebase circuit generates the sampling vector address from a single input clock pin. It also generates internal synchronization signals that govern the conversion. In addition, an independent read system mechanism is available, which minimizes the readout dead time for this particular case of use. The slow control system is composed of a 4 wire SPI like receiver, accepring 1.2 V LVCMOS signals. Its command decoder accepts frames being 64 bit long, containing the targeting internal address space and values to control circuits for various settings and internal biasing voltages (DACs). Figure  2(a) depicts the internal structure of the ASIC and Fig 2b the pictogram of the chip.
With the GRAPH ASIC redesign we address also some desired improvements on the system level [6] , and these are:
@@ 2.2 Sampling windows and memory array
A single sample of the HULA architecture is built using a mixed signal memory, employing a small 30 fC capacitor for analog sampling, a comparator, an overwrite protection flip flop, and a 12 bit digital memory SRAM alike, shown in figure  3(a). Once the sample acquires its analog value, a conversion using the Wilkinson ADC approach will take place at some deterministic point in time. While in conversion, the 12 SRAM bits are connected to the bus and follow the counter value. As soon as the ramp value is above the signal in the capacitor, a D flip flop is set, which disconnects the 12 bit memory from the bus, retaining the last counter value. A separate bus is used to independently read the memory array.
A single sample is stacked into a so-called vertically oriented window containing 64 samples, and there are two banks having 16 windows for each channel. Therefore the memory length of 2048 samples is available for every channel. The memory banks are stacked horizontally, while channels are stacked vertically shown in 3b.
@@ 2.3 Timebase generation circuit
The Timebase generation circuit, figure 4 is composed of a self looped serial shift register being 64 bits in length, inside which a sample actuation signal of programmable width is let to circulate in a loop. When the external nCLR signal is released, the shift of this signal occurs driving the sampling action through the samples in the window at an externally provided clock. At each turnaround, a 5 bit counter increments by one, and its value represents the window which is addressed for sampling, figure  5. The signal from the shift register is directly fanned out to all windows arrays, while the decoded value in the 5 bit counter selects the window undergoing sampling  3(a). In addition, the Timebase extrapolates a few more signals necessary for internal synchronization. An AnB signal is sent off the chip, when the transition from one bank to another happens and internal signals are set to control the conversion controller.
Figure 4: Functional block diagram of the Timebase circuit.
Figure 5: Sampling signal sequence diagram.
@@ 2.4 Conversion controller
The main function of the conversion controller is to execute the conversion at a deterministic moment. It is composed of ramp generator, which relies on a programmable current source driven into a capacitor to create a voltage ramp, figure  6(a). A gray coded counter value is derived from a synchronous counter that revolves at the externally provided frequency clock. The 12 bit Gray code and analog ramp signal are further buffered for fanout over the 16 channels of the entire bank  6. Hence the ASIC has two such controllers, one for each bank.
@@ 2.5 HULA operation
HULA operates in a continuously revolving pattern. Once the sampling starts, the analog signal is sampled in the first Bank. As soon as the first Bank A has completed sampling, the AnB signal changes, signaling the transition to the external world, and automatically starts to sample bank B and puts the entire Bank A into conversion. When the Bank B has completed sampling, Bank A contains digitized data from the previous revolution which are available for the duration of new sampling period of Bank A, figure  7 .
Considering there are 1024 samples in each Bank, selecting a sampling clock of 125 MHz will require 8,192 $\mu$s to complete the sampling of one Bank. To reach 12 bit resolution in this time, the external clock should run at least at 2 $n$s period, hence at 500 MHz, and the ramp should be adjusted to cover the operating headroom. This allows to set an arbitrary sampling frequency, if a lower sampling rate is more adequate for a larger CSA shaping time. It also enables reading a much larger portion of the memory in the given conversion time slot, covering more events.
@@ 2.6 Data readout
The mechanism for memory data readout is completely independent from sampling and conversion, and allows some flexibility in accessing values from the desired address. It is possible to read consecutive samples from the chosen address, while loading a new one, reducing dead time. This is achieved having a receiver composed of 3 serial inputs (channel,window, sample), a corresponding read clock and load confirmation. When an address is shifted in the system and loaded, the value is an offset added to a counter, which is driven by the read clock. At the rising edge of a clock cycle a new vector is decoded pointing to a single cell. This gets attached to the read bus and streamed to a 12 bit parallel output of the chip, which an FPGA latches at the opposite clock edge, figure  8a. Introducing a value of zero to all three registers, and continuing clocking, one would read sample 0 from channel 0, till the last sample of the channel 0. By continuing asserting the read clock, sample 0 of channel 1 would follow and so on. In our application, we load the address of interest, and make the next 6 clock cycles to load new address, while collecting requested data from the previous call. Such access is envisioned for pulse shape extraction. The maximum transmission frequency tested so far was 60 MHz, which would drain the entire memory in loop mode in about 0.5 $m$s, collecting some 50kB of data.
@@ 2.7 Trigger section
Unlike operating in loop mode, in continuous mode triggers are key for the selection of events to pull from GRAPH memory. Each channel has a trigger circuit composed of a programmable threshold, a logical selection for firing on rising or falling edge of the input signal, and a self reset logical signal (the actual trigger output), which pulse width can be set between a ns to few hundred ns by an internal DAC, figure 9(a). Adjacent triggers are OR-ed and fanned out toward the FPGA via an LVDS output 9b. To a degree it allows to discriminate the pulse hit, by setting a common threshold, while setting the width differently. This configuration simplifies the required firmware for channel spatial charge distribution search. Ideally, the FPGA is only required to have an 11 bit counter (2048 sample space) being released by the nCLR signal as for all chips, and increments synchronously with the sampling clock. At trigger arrival the value is latched, and the readout processing sequence can estimate the events of interest to extrapolate the charge cloud signal form the MCP/anode. An additional possibility to enhance the functionality is by adding a TDC on the muxed triggers, and locate with better timming the absolute time of the photon arrival with sub $n$s time of arrival precision should be possible. An additional option, the software trigger, is also available, and can be initiated via the slow control serial register. Both trigger options would stop the ASIC in loop mode operation, as well as, both can be masked to be overruled.
@@ 3 Results
For testing the ASICs, a PCB board that can be attached to a detector was produced, figure 10. This solution has currently 4 ASICs (2 per axis) wire bonded directly onto the PCB along with a minimal set of low drop regulators and a subsets of dual inline connectors to connect to an external FPGA board. A perpendicular connector made of two dual inline pin headers is used to couple the readout system to the detector. For testing, the sampling clock was set to 125MHz, and the Wilkinson conversion clock was set to 500 MHz.
Figure 10: Detector test board
Initially, the CSA amplifier was skipped, and only the performance of the HULA core was examined by carefully tuning the conversion slope to match the pedestal offset on both Banks. Because the comparators in each memory cell cannot operate near the voltage supply rails, using a DC swap the usable ADC headroom was extracted. Although each ASIC is slightly different, we found all of them operate safely within a 800 mV headroom covered by some 3000 ADC counts. The INL has found to be of few percent, figures 11a,11b.
By applying a DC voltage of about 600 mV to the input, the recorded pedestal spread was found to be considerable, with a sigma of 142 counts, shown in figure 12a This somewhat expected result comes from a combination of factors. These are the systematic production mismatch of the capacitors used in the sampling cell, the intrinsic firing point error of the comparator, and the location of cell itself inside the large memory array, as the ramp propagates through the fanout. A solid calibration method is to measure each cell about a 100 times and average its value. As a new measurement is taken, the averaged offset value is subtracted from the measurement. After doing so, the pedestal value is removed, the intrinsic offset of the cell is corrected reducing the spread to a sigma of 23 counts, figure 12b. The ADC has an SNR of 20log*(3000/23) of around 42dB, with the sensitivity of about 6 mV. Although the system is set to use 12 bits of data, the ENOB figure is more around 7. The numbers obtained from the fit are spoiled due to large tails (figure 12b) produced by a small fraction of large values outliers. In time domain these represent single samples that register a large value, and doesn’t fire a trigger, so it seems are more related errors in conversion. The effect isn’t statically related to particular cells but rather happens sporadically. We are sill investigating the cause. By neglecting these single sample off errors, the sigma reduces toward 10 counts.
@@
The system was connected to the detector, and the CSAv3 amplifiers were turned on having the calibration procedure being repeated. The amplifier noise generated by the 5 $p$F input capacitance has worsen the result to 38 ADC counts, figure 13a. It was found that the noise contribution could be reduced by off line post processing the data using a combination of a median and FIR filter to some degree of success without degrading the signal shape, figure 13b. None the less, the over all obtained raw results are close to the specifications and expected performance.
@@ 4 Summary and future work
We developed the GRAPH ASIC, which enables the construction of a readout system intended to couple to a large area MCP single photon imaging detector, designated for future space applications. The test board was constructed and the ASIC performance evaluated. Furthermore initial test confirm the correct functionality as well as correct operation. In addition tests recording detector signals were carried out. Considering the low power consumption of 47 mW per channel, an extrapolated value for a 8x8cm detector requiring 128 channels per axis *(16 ASICs) would yield a power consumption of 13 Watts FPGA excluded. In the future we look forward to testing a new kind of firmware, that would only record events, by asserting a few clock cycles only when a signal is detected via triggers, and force conversion, if an arbitrary overdue time has passed to avoid losing the signal due to leakage. In this way, we aim to optimize the usage of the memory while preserving the sampling clock. Ultimately, an improved version of HULA considering a much better SNR, could operate as an sample and hold event sampler with the ability to digitize only events. In addition, an internal precise TDC with memory could adjunct the time of arrival to the events.